OpenStudy (anonymous):

help will medal what is the slope and y-intercept of this line? y = -6x + 2 A. slope: -6, y-intercept: 2 B. slope: -3, y-intercept: 1 C. slope: 2, y-intercept: -6 D. slope: 6, y-intercept: -2

rishavraj (rishavraj):

y = mx+c m - slope c- y intercept

rishavraj (rishavraj):

coefficient of y should be 1

OpenStudy (priyar):

so here what is the slope? @Mavckenzie ?

OpenStudy (anonymous):

wait what does the m stand for?

rishavraj (rishavraj):

m is for slope

OpenStudy (priyar):

@Mavckenzie which part did u not understand?

OpenStudy (anonymous):

all of it sorry, is there another way to describe the formula?

OpenStudy (priyar):

y=mx+c represents the equation of a line..where 'm' is just a letter used to represent the slope of the line..and 'c' is another letter used to represent the y-intercept.. i have used to terms "slope " and 'y-intercept' do u know these terms?

OpenStudy (priyar):

*two terms

OpenStudy (priyar):

@Mavckenzie ??

OpenStudy (anonymous):

so if they are just numbers to help represent the slope and y- intercept how am I going to use them?

OpenStudy (priyar):

compare the given eq with the general one i have written and find m and c ..u can do so just by looking.

OpenStudy (priyar):

what do u get?

OpenStudy (skullpatrol):

$$\Huge y=mx+c$$ $$\Huge y=–6x+2$$

OpenStudy (priyar):

thanks @skullpatrol ..i think now it would be easier for @Mavckenzie ..

OpenStudy (skullpatrol):

m = ? and c = ?

OpenStudy (anonymous):

i might be wrong but m and c together equals -6 yes? so do I have to just figure out what x is or do I still have to find out what m and c is?

OpenStudy (priyar):

how is m and c together -6?

OpenStudy (priyar):

u must compare the coefficient of x separately...

OpenStudy (anonymous):

B?

OpenStudy (priyar):

No..try again

OpenStudy (skullpatrol):

There is NOTHING to "figure out"... Just look under the m and tell me the number Look under the c and tell me the number.

OpenStudy (priyar):



OpenStudy (priyar):

look at the arrows (lines) and tell again...

OpenStudy (priyar):

m= ?

OpenStudy (skullpatrol):

m = slope = ?

OpenStudy (skullpatrol):

c = y-intercept = ?

OpenStudy (priyar):

do u need an example? @Mavckenzie ?

OpenStudy (skullpatrol):

The "?" Means what number goes here :-)

OpenStudy (priyar):

wait..let me give u a similar example.. if they had given y=3x+2 i would have said m=3 and c=2..by comparing the eq.. with y=mx +c similarly for this eq what will be your m and c?

OpenStudy (skullpatrol):

$$\Huge y = (slope)x+(y-intercept)$$ $$\Huge y = (m)x + c$$ $$\Huge y=(-6)x+ 2$$

OpenStudy (skullpatrol):

slope = m = ? and y-intercept = c = ?

OpenStudy (anonymous):

i'm pretty sure it is b

OpenStudy (anonymous):

For the equation y = mx + b, m is equal to the slope and b is equal to the y-intercept.

OpenStudy (skullpatrol):

Why are you sure that the slope = -3 and the y-intercept = 1 ?

OpenStudy (anonymous):

its 1am and I already submitted my answer and I got it wrong because I was told b was wrong but it is telling me For the equation y = mx + b, m is equal to the slope and b is equal to the y-intercept.

OpenStudy (anonymous):

ok I already submitted my answers at this time, I know c is wrong so what do you think it is?

OpenStudy (skullpatrol):

what is the slope and y-intercept of this line? y = -6x + 2 <----> y = mx + b A. slope: -6, y-intercept: 2 <--------------------------^----^-------------^-----^ B. slope: -3, y-intercept: 1 C. slope: 2, y-intercept: -6 D. slope: 6, y-intercept: -2